# On cutoff via rigidity for high dimensional curved diffusions

### v3

@@ 2.1. Fisher information and the heat kernel
It is natural to try to directly upper bound the Fisher information by using the heat kernel, instead of using a comparison with relative entropy. Let us examine first the exactly solvable case of the OU process, for which the Mehler formula $\mu_{t}=\mathcal{N}(\mathrm{e}^{-\rho t}x_{0},(1-\mathrm{e}^{-2\rho t})\mathrm{I}_{d})$ gives, for $t>0$,
$\nabla\log\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\mu}(x)=-\frac{x-\mathrm{e}^{-\rho t}x_{0}}{1-\mathrm{e}^{-2\rho t}}+x=a_{t}x+b_{t}x_{0}$ (2.1)
@@
$\mathrm{I}(\mu_{t_{0}}\mid\mu)\leq 2a_{t_{0}}^{2}\int|x|^{2}\mathrm{d}\mu(x)+2b_{t_{0}}^{2}|x_{0}|^{2}=2a_{t_{0}}^{2}d+2b_{t_{0}}^{2}|x_{0}|^{2}.$ (2.2)
This gives the upper bound (3.32). Beyond the OU process, denoting, for all $t>0$, by $p_{t,x_{0}}$ the density of $\mu_{t}$ with respect to the Lebesgue measure, which is known as the heat kernel of $\mathcal{L}$, we have, for all $x$ and all $t>0$,
@@
$\bigr|\nabla\log p_{t,x_{0}}(x)\bigr|^{2}\leq a_{t}^{2}|x|^{2}+b_{t}^{2}|x_{0}|^{2},$ (2.4)
@@
$\mathrm{I}(\mu_{t_{0}}\mid\mu)\leq 2\int(a_{t_{0}}^{2}|x|^{2}+|\nabla V(x)|^{2})\mathrm{d}\mu_{t}(x)+2b_{t_{0}}^{2}|x_{0}|^{2}.$ (2.5)
Here $a_{t}$ and $b_{t}$ are constants with respect to $x$, that may depend on $t$ but not on $d$ and $x_{0}$. The explicit dependence over the dimension and initial condition is crucial here for our purposes. Unfortunately, we were not able to locate (2.4) in the literature on short time logarithmic gradient bounds for the heat kernel on non-compact manifolds with a lower curvature bound. Close works are Engoulatov (2006) where a constant is not controlled with respect to the dimension, and Chen et al. (2023) which is curvature free but is restricted to a compact subset.
To summarize, the upper bound on $\mathrm{I}(\mu_{t}\mid\mu)$ follows from a logarithmic gradient bound for the heat kernel in short time, which expresses a regularization property of the evolution equation driven by $\mathcal{L}$ when its starts from a Dirac mass. Such an upper bound on the Fisher information $\mathrm{I}$ implies all the others. Indeed, it would give an upper bound on $\mathrm{H}$ by using the logarithmic Sobolev inequality, giving in turn an upper bound on $\mathrm{W}_{2}$ by using the Talagrand inequality, as well an upper bound on $\mathrm{d}_{\mathrm{TV}}$ by using the Pinsker inequality.
@@ 2.3. Normalization
For any real $\alpha>0$ which may depend on $d$, the time-changed process $X^{(\alpha)}:={(X_{\alpha t})}_{t\geq 0}$ solves $\mathrm{d}X^{(\alpha)}_{t}=-\alpha\nabla V(X^{(\alpha)}_{t})\mathrm{d}t+\sqrt{2\alpha}\mathrm{d}B_{t}$ and has generator $\alpha\mathcal{L}$. The process $X^{(\alpha)}$ has cutoff at critical time $t_{*}$ iif $X^{(\alpha)}$ has cutoff at critical time $t_{*}/\alpha$. Such scaled processes play a role with respect to mean-field limits of interacting particle systems related to McKean – Vlasov semilinear PDE, see Boursier et al. (2023). Note that the product condition is invariant under such a scaling.
@@ 3.1. Preliminaries on rigidity
As already mentioned, our starting point is a Gaussian splitting or factorization theorem of Cheng and Zhou (2017), which in the Euclidean setting takes the form of Theorem 3.1 below. See also (Chafaï and Lehec, 2020, Rem. 3.2).
@@
$V(x)=\frac{\rho}{2}((x_{1}-m_{1})^{2}+\cdots+(x_{k}-m_{k_{1}})^{2})+\widetilde{V}(x_{k_{1}+1},...,x_{d})$ (3.1)
where $k_{1}$ is the dimension of the eigenspace associated with $\lambda_{1}$, and $\widetilde{V}$ is $\rho$-convex on $\mathbb{R}^{d-k_{1}}$. Moreover, all eigenfunctions with eigenvalue $\lambda_{1}$ are affine, and only depend on the first $k_{1}$ coordinates in the above basis.
@@
Since eigenfunctions satisfy
$\int|\nabla f|^{2}\mathrm{d}\mu=-\int(\mathcal{L}f)f\mathrm{d}\mu=\lambda_{1}\int f^{2}\mathrm{d}\mu,$ (3.2)
any eigenfunction with eigenvalue $\lambda_{1}$ is of the form $\langle a,p(x)\rangle_{\mathbb{R}^{k}}+b$ with $p$ the projection on the $k$-dimensional Gaussian factor and $|a|^{2}=\lambda_{1}\|f\|_{L^{2}(\mu)}^{2}$.
Another way of stating this result is that up to a rotation (the change of basis) and a translation by the vector $m$, the law $\mathrm{e}^{-V}$ is a product measure, with a centered Gaussian factor with variance $\rho$ on the first $k$ coordinates, and a log-concave factor with a $\rho$-convex potential on the last dkd-k coordinates. Note that the result can only be true up to a rotation and translation, since the assumptions are stable by isometries of $\mathbb{R}^{d}$.
@@
We shall show that $T(X)$ is a Markov process and recognize it as an OU process by computing the generator. If $\vec{v}$ is a vector-valued function, then $\vec{\mathcal{L}}\vec{v}$ is the vector obtained by applying $\mathcal{L}$ to each coordinate, and $\Gamma(\vec{v})$ the matrix whose coefficients are $\Gamma(\vec{v}_{i},\vec{v}_{j})$. Since the coordinates of $T$ are orthogonal normalized eigenfunctions, $\vec{\mathcal{L}}T=-\lambda_{1}T$ and $\Gamma(T)=\rho\operatorname{Id}$, as per Remark 3.2. By the diffusion property, we have
@@
This is a function of $T$, so $T(X)$ is a Markov process, and when viewing it as such it is indeed the generator of an OU process with variance $\rho^{-1}$, applied to a function $g$. ∎
@@
Let us now prove the lower bound. Consider a multi-eigenmap $T$, see Definition 1.8. Up to an isometry, it is of the form $T(x)=Ax+b$. Moreover, since the coordinates of $T$ are orthogonal eigenfunctions, the columns of $A$ are orthogonal vectors. If moreover we assume that $T\in\mathbb{S}_{F_{1}}$, then each column of $A$ has norm $\sqrt{\rho}=\sqrt{\lambda_{1}}$. Hence we get
@@
Applying Lemma 3.3, $Z_{t}:=T(X_{t})$ and $T(Y_{t})$ are OU processes with speed accelerated by a factor $\rho$. Suppose now that $Y_{0}$ is distributed according to the equilibrium measure : $Y_{0}\sim\mu$. Then $T(Y_{t})$ follows a standard Gaussian law $\gamma_{k_{1}}$ on $\mathbb{R}^{k_{1}}$, for all $t$. Thus using the formula $\mathrm{W}_{2}(\mathcal{N}(m_{1},\Sigma_{1}),\mathcal{N}(m_{2},\Sigma_{2}))^{2}=(m_{1}-m_{2})^{2}+\mathrm{Tr}((\sqrt{\Sigma_{1}}-\sqrt{\Sigma_{2}})^{2})$, the Mehler formula $Z_{t}\sim\mathcal{N}(z_{0}\mathrm{e}^{-\rho t},(1-\mathrm{e}^{-2\rho t})I_{k_{1}})$, and $\gamma_{k_{1}}=\mathcal{N}(0,I_{k_{1}})$, we find
$\mathrm{W}_{2}(Z_{t},\gamma_{k_{1}})^{2}=\mathrm{e}^{-2\rho t}(|z_{0}|^{2}+k_{1}).$ (3.8)
Hence
$\mathrm{W}_{2}(X_{t},\mu)\geq\frac{\mathrm{e}^{-\rho t}}{\sqrt{\rho}}\sqrt{|T(x_{0})|^{2}+k_{1}}.$ (3.9)
@@
We now consider the case where $\mu$ is centered and $S=B(0,R)$. For any $T\in\mathbb{S}_{F_{1}}$, writing $T=Ax+b$, since $\int T\mathrm{d}\mu=0$ we see that $b=0$. Therefore $T(0)=0$. Moreover, since the columns of $A$ are orthogonal and have norm $\sqrt{\lambda_{1}}$, we get $\sup_{x\in S}|Ax|^{2}=\rho R^{2}$.
@@
Without loss of generality, we can assume that $m^{(d)}=0$ by translating $V^{(d)}$ We start by using the lower bound of Theorem 1.1 to prove the convergence to infinity when $t_{d}=(1-\epsilon)t_{*}$. Since we are in the centered setting and the set of initial conditions is a centered ball of radius $c\sqrt{d}$, the lower bound is
$\frac{\mathrm{e}^{-\rho t}}{\sqrt{\rho}}(\rho c^{2}d+k_{1})^{1/2}\leq\mathrm{W}_{2}(X_{t},\mu).$ (3.10)
Evaluating at $t=t_{d}$ and neglecting $k_{1}$, we get
$cd^{\epsilon/2}\leq\mathrm{W}_{2}(X_{t_{d}},\mu)$ (3.11)
and letting $d$ go to infinity concludes the proof of the lower bound. Let us now prove the case $t_{d}=(1+\epsilon)t_{*}$ via the upper bound in Theorem 1.1. Since $\mu^{(d)}$ is centered,
@@
$\int f^{2}\mathrm{d}\mu\leq\frac{1}{\lambda_{1}^{(d)}}\int|\nabla f|^{2}\mathrm{d}\mu.$ (3.13)
@@
$\int|x|^{2}\mathrm{d}\mu\leq\frac{d}{\lambda_{1}^{(d)}}.$ (3.14)
@@
$\mathrm{W}_{2}(X_{t},\mu)\leq\mathrm{e}^{-\lambda_{1}^{(d)}t}\Bigr(\sup_{x_{0}\in S}|x_{0}|^{2}+\frac{d}{\lambda_{1}^{(d)}}\Bigr)^{1/2}.$ (3.15)
@@
$\mathrm{W}_{2}(X_{t_{d}},\mu)\leq d^{-\epsilon/2}(c^{2}+(\lambda_{1}^{(d)})^{-1})^{1/2}.$ (3.16)
Since $\liminf_{d\to\infty}\lambda_{1}^{(d)}>0$, letting $d$ go to infinity concludes the proof.
Remark 3.4.
In the proof, the assumption $\liminf_{d\to\infty}\lambda_{1}^{(d)}>0$ could have been replaced by a slow enough growth of $(\lambda_{1}^{(d)})^{-1}$, for instance $(\lambda_{1}^{(d)})^{-1}\leq(\log d)^{\alpha}$ for some $\alpha>0$.
@@
Let us prove first the cutoffs in relative entropy and in total variation distance by using Corollary 1.4 on the Wasserstein distance. Let us drop the superscript $(d)$ to simplify the notation.
@@
$\mathrm{H}(X_{t}\mid\mu)\leq\mathrm{e}^{-\rho(t-t_{0})}\mathrm{H}(X_{t_{0}}\mid\mu).$ (3.17)
@@
Using (3.18) with $t=t_{0}=1$, and combining with (3.17), we obtain, for all $t\geq 1$,
$2\mathrm{H}(X_{t}\mid\mu)\leq\mathrm{e}^{-\rho(t-1)}\mathrm{W}_{2}(X_{0},\mu)^{2}.$ (3.19)
@@
$\mathrm{d}_{\mathrm{TV}}(\mu,\nu)^{2}\leq 2\mathrm{H}(\nu\mid\mu),$ (3.20)
@@
$\mathrm{d}_{\mathrm{TV}}(X_{t},\mu)^{2}\leq 2\mathrm{H}(X_{t}\mid\mu)\leq\mathrm{e}^{-\rho(t-1)}\mathrm{W}_{2}(X_{0},\mu)^{2}.$ (3.21)
Now, since $t_{*}\to+\infty$ as $d\to\infty$ in Corollary 1.3, we get, for all $c>0$ and $\varepsilon>0$,
@@
This contractibility argument is also at the heart of the lower bounds in Boursier et al. (2023). It follows that we can bound from below the relative entropy and total variation mixing times by those of a suitable OU process, by taking a multi-eigenfunction map. The lower bounds for OU processes have been established for example in Lachaud (2005), Barrera et al. (2006), and (Boursier et al., 2023, Th. 1.2). As a consequence, we have, in the setting of Corollary 1.4, for all $\varepsilon>0$,
@@
$\mathrm{W}_{2}(X_{t},\mu)^{2}\leq\frac{2}{\rho}\mathrm{H}_{\mu}(X_{t})$ (3.25)
@@
$\mathrm{H}(\nu\mid\mu^{(d)})\leq\frac{1}{2\rho}\mathrm{I}(\nu\mid\mu^{(d)}).$ (3.26)
Taking $\nu=\mathrm{Law}(X^{(d)}_{t})$ and using the cutoff lower bound for $\mathrm{H}$ give, for all $\varepsilon>0$,
$\lim_{d\to\infty}\sup_{x^{(d)}_{0}\in S^{(d)}}\mathrm{I}(X^{(d)}_{t_{d}}\mid\mu^{(d)})=+\infty\quad\text{where}\quad t_{d}=(1-\varepsilon)t_{*}.$ (3.27)
@@
$\partial_{t}\mathrm{H}_{t}(\mu_{t}\mid\mu)=-\mathrm{I}(\mu_{t}\mid\mu)\leq 0\quad\text{and}\quad\partial_{t}\mathrm{I}(\mu_{t}\mid\mu)\leq-2\rho\mathrm{I}(\mu_{t}\mid\mu)\leq 0$ (3.28)
where $\mu_{t}$ and $\mu$ stand for $\mathrm{Law}(X^{(d)}_{t})$ and $\mu^{(d)}$. The Grönwall lemma implies then the exponential decay of Fisher information : for all $t\geq t_{1}>0$,
$\mathrm{I}(\mu_{t}\mid\mu)\leq\mathrm{e}^{-2\rho(t-t_{0})}\mathrm{I}(\mu_{t_{1}}\mid\mu)$ (3.29)
@@
where we have used the monotinicity of $\mathrm{I}$. This gives, combined with (3.19), when $t_{0}>1$,
$\mathrm{I}(\mu_{t_{1}}\mid\mu)\leq\frac{\mathrm{H}(\mu_{t_{0}}\mid\mu)}{t_{1}-t_{0}}\leq\frac{\mathrm{e}^{-\rho(t_{0}-1)}}{2(t_{1}-t_{0})}\mathrm{W}_{2}(X_{0},\mu)^{2}.$ (3.31)
This gives finally a cutoff upper bound : for all $\varepsilon>0$,
$\lim_{d\to\infty}\sup_{x^{(d)}_{0}\in S^{(d)}}\mathrm{I}(X^{(d)}_{t_{d}}\mid\mu^{(d)})=0\quad\text{where}\quad t_{d}=(1+\varepsilon)t_{*}.$ (3.32)
@@
Let $h(x)=x_{1}+\cdots+x_{d}$ be the symmetric Hermite polynomial of first degree. Its gradient is the constant vector $\nabla h=(1,\ldots,1)$. Since (1.5) holds, we have $\lambda_{1}\geq\rho$. Hence, to get $\lambda_{1}=\rho$, it suffices to show that $h$ is an eigenfunction of -$\mathcal{L}$ with eigenvalue $\rho$, and then apply the splitting theorem. The generator is

### v4

@@ 2.1. Fisher information and the heat kernel
It is natural to try to directly upper bound the Fisher information by using the heat kernel, instead of using a comparison with relative entropy. Following Boursier et al. (2023), let us examine first the exactly solvable case of the OU process, for which the Mehler formula $\mu_{t}=\mathcal{N}(\mathrm{e}^{-\rho t}x_{0},\frac{1-\mathrm{e}^{-2\rho t}}{\rho}\mathrm{I}_{d})$ gives, for $t>0$,
$\nabla\log\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\mu}(x)=-\rho\frac{x-\mathrm{e}^{-\rho t}x_{0}}{1-\mathrm{e}^{-2\rho t}}+\rho x=a_{t}x+b_{t}x_{0}$ (2.1)
@@
$\mathrm{I}(\mu_{t}\mid\mu)=\rho d\frac{\mathrm{e}^{-4\rho t}}{1-\mathrm{e}^{-2\rho t}}+\rho^{2}\mathrm{e}^{-2\rho t}|x_{0}|^{2}.$ (2.2)
This gives the upper bound (3.32) as well as the lower bound (3.27). Beyond the OU process, denoting, for all $t>0$, by $p_{t,x_{0}}$ the density of $\mu_{t}$ with respect to the Lebesgue measure, which is known as the heat kernel of $\mathcal{L}$, we have, for all $x$ and all $t>0$,
@@
$\bigl|\nabla\log p_{t,x_{0}}(x)\bigr|^{2}\leq a_{t}^{2}|x|^{2}+b_{t}^{2}|x_{0}|^{2},$ (2.4)
@@
$\mathrm{I}(\mu_{t}\mid\mu)\leq 2\int(a_{t}^{2}|x|^{2}+|\nabla V(x)|^{2})\mathrm{d}\mu_{t}(x)+2b_{t}^{2}|x_{0}|^{2}.$ (2.5)
Here $a_{t}$ and $b_{t}$ are constants with respect to $x$, that may depend on $t$ but not on $d$ and $x_{0}$. The explicit dependence on the dimension and initial condition is crucial here for our purposes. Unfortunately, we were not able to locate (2.4) in the literature on short time logarithmic gradient bounds for the heat kernel on non-compact manifolds with a lower curvature bound. Close works are Engoulatov (2006) where a constant is not controlled with respect to the dimension, and Chen et al. (2023) which is curvature free but is restricted to a compact subset. It would not be a surprise to assume, in addition to positive curvature, an additional condition regarding eigenfunctions associated to the spectral gap.
To summarize, the upper bound on $\mathrm{I}(\mu_{t}\mid\mu)$ follows from a logarithmic gradient bound for the heat kernel in short time, which expresses a regularization property of the evolution equation driven by $\mathcal{L}$ when it starts from a Dirac mass. Such an upper bound on the Fisher information $\mathrm{I}$ implies all the others. Indeed, it would give an upper bound on $\mathrm{H}$ by using the logarithmic Sobolev inequality, giving in turn an upper bound on $\mathrm{W}_{2}$ by using the Talagrand inequality, as well as an upper bound on $\mathrm{d}_{\mathrm{TV}}$ by using the Pinsker inequality.
@@ 2.3. Normalization
For any real $\alpha>0$ which may depend on $d$, the time-changed process $X^{(\alpha)}:={(X_{\alpha t})}_{t\geq 0}$ solves $\mathrm{d}X^{(\alpha)}_{t}=-\alpha\nabla V(X^{(\alpha)}_{t})\mathrm{d}t+\sqrt{2\alpha}\mathrm{d}B_{t}$ and has generator $\alpha\mathcal{L}$. The process $X$ has cutoff at critical time $t_{*}$ iff $X^{(\alpha)}$ has cutoff at critical time $t_{*}/\alpha$. Such scaled processes play a role with respect to mean-field limits of interacting particle systems related to McKean–Vlasov semilinear PDE, see Boursier et al. (2023). Note that the product condition is invariant under such a scaling.
@@ 3.1. Preliminaries on rigidity
As already mentioned, our starting point is a Gaussian splitting or factorization theorem of Cheng and Zhou (2017), which in the Euclidean setting takes the form of Theorem 3.1 below. See also Section 1.5 and (Chafaï and Lehec, 2020, Rem. 3.2).
@@
$V(x)=\frac{\rho}{2}((x_{1}-m_{1})^{2}+\cdots+(x_{k_{1}}-m_{k_{1}})^{2})+\widetilde{V}(x_{k_{1}+1},\ldots,x_{d})$ (3.1)
where $k_{1}$ is the dimension of the eigenspace $E_{1}$ of -$\mathcal{L}$ associated with $\lambda_{1}$, and $\widetilde{V}$ is $\rho$-convex on $\mathbb{R}^{d-k_{1}}$. Moreover, all eigenfunctions with eigenvalue $\lambda_{1}$ are affine, and only depend on the first $k_{1}$ coordinates in the above basis.
@@
By Theorem 3.1, if $f\in E_{1}$, then $f$ is affine, and since $\int|\nabla f|^{2}\mathrm{d}\mu=-\int(\mathcal{L}f)f\mathrm{d}\mu=\lambda_{1}\int f^{2}\mathrm{d}\mu$, we get $f(x)=\langle a,p(x)\rangle_{\mathbb{R}^{k_{1}}}+b$ with $p$ the projection on the $k_{1}$-dimensional Gaussian factor and $|a|^{2}=\lambda_{1}\|f\|_{L^{2}(\mu)}^{2}$.
Another way of stating this result is that up to a rotation (the change of basis) and a translation by the vector $m$, the law $\mathrm{e}^{-V}$ is a product measure, with a centered Gaussian factor with variance $1/\rho$ on the first $k_{1}$ coordinates, and a log-concave factor with a $\rho$-convex potential on the last $d-k_{1}$ coordinates. Note that the result can only be true up to a rotation and translation, since the assumptions are stable by isometries of $\mathbb{R}^{d}$.
@@
Let us show that $T(X)$ is a Markov process and compute its generator. If $\vec{v}$ is a vector-valued function, then $\vec{\mathcal{L}}\vec{v}$ is the vector obtained by applying $\mathcal{L}$ to each coordinate, and $\Gamma(\vec{v})$ the matrix whose coefficients are $\Gamma(\vec{v}_{i},\vec{v}_{j})$. Since the coordinates of $T$ are orthogonal normalized eigenfunctions, $\vec{\mathcal{L}}T=-\lambda_{1}T$ and $\Gamma(T)=\rho\operatorname{Id}$, as per Theorem 3.1 (affine with orthogonal constant gradients of norm $\sqrt{\rho}$). By the diffusion property, we have
@@
This is a function of $T$, so $T(X)$ is a Markov process, and when viewing it as such it is indeed equal to $\rho$ times the generator of an OU process with unit variance, applied to a function $g$. ∎
@@
Let us now prove the lower bound. Consider a multi-eigenmap $T$, see Definition 1.8. Up to an isometry, it is of the form $T(x)=Ax+b$. Moreover, since the coordinates of $T$ are orthogonal eigenfunctions, the rows of $A$ are orthogonal vectors. If moreover we assume that $T\in\mathbb{S}_{F_{1}}$, then each row of $A$ has norm $\sqrt{\rho}=\sqrt{\lambda_{1}}$. Hence we get
@@
Applying Lemma 3.3, $Z_{t}:=T(X_{t})$ and $T(Y_{t})$ are OU processes with speed accelerated by a factor $\rho$. Suppose now that $Y_{0}$ is distributed according to the equilibrium measure: $Y_{0}\sim\mu$. Then $T(Y_{t})$ follows a standard Gaussian law $\gamma_{k_{1}}$ on $\mathbb{R}^{k_{1}}$, for all $t$. Thus using the formula $\mathrm{W}_{2}(\mathcal{N}(m_{1},\Sigma_{1}),\mathcal{N}(m_{2},\Sigma_{2}))^{2}=|m_{1}-m_{2}|^{2}+\mathrm{Tr}((\sqrt{\Sigma_{1}}-\sqrt{\Sigma_{2}})^{2})$ when $\Sigma_{1}\Sigma_{2}=\Sigma_{2}\Sigma_{1}$, the Mehler formula $Z_{t}\sim\mathcal{N}(T(x_{0})\mathrm{e}^{-\rho t},(1-\mathrm{e}^{-2\rho t})I_{k_{1}})$, and $\gamma_{k_{1}}=\mathcal{N}(0,I_{k_{1}})$, we find
$\mathrm{W}_{2}(Z_{t},\gamma_{k_{1}})^{2}=\mathrm{e}^{-2\rho t}|T(x_{0})|^{2}+k_{1}\bigl(1-\sqrt{1-\mathrm{e}^{-2\rho t}}\bigr)^{2}.$ (3.7)
@@
We now consider the case where $\mu$ is centered and $S=B(0,R)$. For any $T\in\mathbb{S}_{F_{1}}$, writing $T=Ax+b$, since $\int T\mathrm{d}\mu=0$ we see that $b=0$. Therefore $T(0)=0$. Moreover, since the rows of $A$ are orthogonal and have norm $\sqrt{\lambda_{1}}$, we get $\sup_{x\in S}|Ax|^{2}=\lambda_{1}R^{2}$.
@@
Let us drop the $(d)$ superscript. Without loss of generality, we can assume that $m=0$ by translating $V$. We start by using the lower bound of Theorem 1.1 to prove the convergence to infinity when $t_{d}=(1-\epsilon)t_{*}$. Since we are in the centered setting and the set of initial conditions is a centered ball of radius $c\sqrt{d}$, the lower bound
$\frac{\mathrm{e}^{-2\lambda_{1}t}}{\lambda_{1}}\sup_{x_{0}\in S}\Lambda(x_{0})\leq\sup_{x_{0}\in S}\mathrm{W}_{2}(X_{t},\mu)^{2}$ (3.8)
while $\sup_{x_{0}\in S}\Lambda(x_{0})=\lambda_{1}c^{2}d$ thanks to Theorem 1.1 since $S=B(0,c\sqrt{d})$ with $c>0$. Evaluating at $t=t_{d}$ and neglecting $k_{1}$, we get
$cd^{\epsilon/2}\leq\sup_{x_{0}\in S}\mathrm{W}_{2}(X_{t_{d}},\mu)$ (3.9)
and letting $d$ go to infinity concludes the proof of the lower bound. Let us now prove the case $t_{d}=(1+\epsilon)t_{*}$ via the upper bound in Theorem 1.1. Since $\mu$ is centered,
@@
$\int f^{2}\mathrm{d}\mu\leq\frac{1}{\lambda_{1}}\int|\nabla f|^{2}\mathrm{d}\mu.$ (3.11)
@@
$\int|x|^{2}\mathrm{d}\mu\leq\frac{d}{\lambda_{1}}.$ (3.12)
@@
$\sup_{x_{0}\in S}\mathrm{W}_{2}(X_{t},\mu)\leq\mathrm{e}^{-\lambda_{1}t}\Bigl(\sup_{x_{0}\in S}|x_{0}|^{2}+\frac{d}{\lambda_{1}}\Bigr)^{1/2}.$ (3.13)
@@
$\sup_{x_{0}\in S}\mathrm{W}_{2}(X_{t_{d}},\mu)\leq d^{-\epsilon/2}(c^{2}+(\lambda_{1})^{-1})^{1/2}.$ (3.14)
Since $\liminf_{d\to\infty}\lambda_{1}>0$, letting $d$ go to infinity concludes the proof.
Remark 3.4 (Weaker assumption).
In the proof, the assumption $\liminf_{d\to\infty}\lambda_{1}>0$ could have been replaced by a slow enough growth of $(\lambda_{1})^{-1}$, for instance $(\lambda_{1})^{-1}\leq(\log d)^{\alpha}$ for some $\alpha>0$.
@@
Let us prove first the cutoffs in relative entropy and in total variation distance by using Corollary 1.3 on the Wasserstein distance. We drop the $(d)$ superscript.
@@
$\mathrm{H}(X_{t}\mid\mu)\leq\mathrm{e}^{-2\rho(t-t_{0})}\mathrm{H}(X_{t_{0}}\mid\mu).$ (3.15)
@@
Using (3.16) with $t=1$, and combining with (3.15), we obtain, for all $t\geq 1$,
$2\mathrm{H}(X_{t}\mid\mu)\leq\mathrm{e}^{-2\rho(t-1)}\mathrm{W}_{2}(X_{0},\mu)^{2}.$ (3.17)
@@
$\mathrm{d}_{\mathrm{TV}}(\mu,\nu)^{2}\leq\frac{1}{2}\mathrm{H}(\nu\mid\mu),$ (3.18)
@@
$\mathrm{d}_{\mathrm{TV}}(X_{t},\mu)^{2}\leq\frac{1}{2}\mathrm{H}(X_{t}\mid\mu)\leq\frac{1}{4}\mathrm{e}^{-2\rho(t-1)}\mathrm{W}_{2}(X_{0},\mu)^{2}.$ (3.19)
Recall that $\rho=\lambda_{1}$. Now, proceeding as in the proof of Corollary 1.3, we get, for all $c>0$,
$\sup_{x_{0}\in B(m,c\sqrt{d})}\mathrm{W}_{2}(X_{0},\mu)^{2}=\sup_{x_{0}\in B(m,c\sqrt{d})}\int|x-x_{0}|^{2}\mathrm{d}\mu(x)\leq(c^{2}+\lambda_{1}^{-1})d.$ (3.20)
Using the definition of $t_{*}$ and the assumption $\lambda_{1}=o_{d\to\infty}(\log(d))$, we get, for all $c,\varepsilon>0$,
$\sup_{x_{0}\in B(m,c\sqrt{d})}\mathrm{e}^{-2\rho((1+\varepsilon)t_{*}-1)}\mathrm{W}_{2}(X_{0},\mu)^{2}\leq\mathrm{e}^{O(\lambda_{1})}d^{-\varepsilon}\xrightarrow[d\to\infty]{}0.$ (3.21)
Therefore, going back to $\mathrm{H}$ and $\mathrm{d}_{\mathrm{TV}}$, we get, for all $c,\varepsilon>0$,
@@
This contractivity argument is also at the heart of the lower bounds in Boursier et al. (2023). It follows that we can bound from below the relative entropy and total variation mixing times by those of a suitable OU process, by taking a multi-eigenfunction map. The lower bounds for OU processes have been established for example in Lachaud (2005), Barrera et al. (2006), and (Boursier et al., 2023, Th. 1.2). Therefore, in the setting of Corollary 1.4, for all $c>0$, $\varepsilon>0$,
@@
$\mathrm{W}_{2}(X_{t},\mu)^{2}\leq\frac{2}{\rho}\mathrm{H}(X_{t}\mid\mu)$ (3.25)
@@
$\mathrm{H}(\nu\mid\mu)\leq\frac{1}{2\rho}\mathrm{I}(\nu\mid\mu).$ (3.26)
Taking $\nu=\mathrm{Law}(X_{t})$ and using the cutoff lower bound for $\mathrm{H}$ give, for all $\varepsilon>0$,
$\lim_{d\to\infty}\sup_{x_{0}\in S}\mathrm{I}(X_{t_{d}}\mid\mu)=+\infty\quad\text{where}\quad t_{d}=(1-\varepsilon)t_{*}.$ (3.27)
@@
$\partial_{t}\mathrm{H}(\mu_{t}\mid\mu)=-\mathrm{I}(\mu_{t}\mid\mu)\leq 0\quad\text{and}\quad\partial_{t}\mathrm{I}(\mu_{t}\mid\mu)\leq-2\rho\mathrm{I}(\mu_{t}\mid\mu)\leq 0$ (3.28)
where $\mu_{t}$ and $\mu$ stand for $\mathrm{Law}(X_{t})$ and $\mu$. The Grönwall lemma implies then the exponential decay of Fisher information: for all $t\geq t_{1}>0$,
$\mathrm{I}(\mu_{t}\mid\mu)\leq\mathrm{e}^{-2\rho(t-t_{1})}\mathrm{I}(\mu_{t_{1}}\mid\mu)$ (3.29)
@@
where we have used the monotonicity of $\mathrm{I}$. This gives, combined with (3.17), when $t_{0}>1$,
$\mathrm{I}(\mu_{t_{1}}\mid\mu)\leq\frac{\mathrm{H}(\mu_{t_{0}}\mid\mu)}{t_{1}-t_{0}}\leq\frac{\mathrm{e}^{-2\rho(t_{0}-1)}}{2(t_{1}-t_{0})}\mathrm{W}_{2}(X_{0},\mu)^{2}.$ (3.31)
Finally, by using this inequality with $t_{0}=t_{d}-1$ and $t_{1}=t_{d}$ where $t_{d}=(1+\varepsilon)t_{*}$ for an arbitrary fixed $\varepsilon>0$, using (3.20), and $\rho=\lambda_{1}=o_{d\to\infty}(\log(d))$, we get
$\lim_{d\to\infty}\sup_{x_{0}\in S}\mathrm{I}(X_{t_{d}}\mid\mu)=0.$ (3.32)
Remark 3.5 (Weaker assumption).
In the proof, the assumption $\lambda_{1}=o(\log(d))$ could have been replaced by $\lambda_{1}=d^{o(1)}$. This can be done by replacing, in the regularization step, the time $1$ by time $\asymp 1/\lambda_{1}$, which produces a prefactor $\lambda_{1}$ for $\mathrm{H}$, hence $\sqrt{\lambda_{1}}$ for $\mathrm{d}_{\mathrm{TV}}$, and $\lambda_{1}^{2}$ for $\mathrm{I}$.
@@
Let $h(x)=x_{1}+\cdots+x_{d}$ be the first symmetric Hermite polynomial. Its gradient is the constant vector $\nabla h=(1,\ldots,1)$. Since $\mathrm{Hess}(V)=\rho I+\mathrm{Hess}(W)\geq\rho I$ as quadratic forms, we have $\lambda_{1}\geq\rho$. Hence, to get $\lambda_{1}=\rho$, it suffices to show that $h$ is an eigenfunction of -$\mathcal{L}$ with eigenvalue $\rho$, and then apply the splitting theorem. The generator is